# An inhomogeneous adiabatic preparation of a quantum critical ground state in two dimensions

### v1

@@ I Introduction
Figure 1: Inhomogeneous ramp. Here, $\epsilon$ is a dimensionless parameter in a Hamiltonian that measures the distance to the quantum critical point. Panel (a) shows an inhomogeneous ramp in a 1D chain at two different times. The transition begins at the center of the chain, where the critical $\epsilon=0$ is reached first, and from where it expands with velocity $v$ towards the chain’s ends. Panel (b) shows an inhomogeneous ramp in a $10\times 10$ square lattice. The critical region with $\epsilon=0$ expands with velocity $v$ from the center towards the edges of the lattice. Both panels (a) and (b) demonstrate ramps with slope $\alpha=1$.
When the critical $H_{f}$ is approached uniformly in space, the AP ramp is a “half-KZM” ramp, ending at the critical point instead of crossing it, and the KZM power laws still apply, as discussed at length in the following Sec. II. A full KZM ramp can be made adiabatic by making it inhomogeneous in space [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79]. Initially, the parameter $\lambda(t,{\bf r})$ can be driven across the critical point at the center of the lattice, then the central region with $H_{f}$ is expanded towards the lattice edges with a velocity $v$ slower than the speed of sound at the critical point.
An inhomogeneous ramp has also been considered as a means to prepare a critical ground state, either in a 1D chain [80] or a quasi-1D stripe [81]. These studies focused on slightly supersonic ramps that create excitations, which follow the ramp towards the ends of the chain, leaving behind the critical ground state in the bulk. Since the excitations eventually bounce back from the ends, we prefer to focus on subsonic ramps in this work. Furthermore, instead of using excitation energy as a cost function—which under-represents low-frequency excitations that can distort the target state without significantly increasing its energy—we focus on the density of excited quasiparticles (for integrable systems) or, more generally, on the (in)fidelity between the critical ground state and the final state (for non-integrable systems).
@@ II A uniform ramp to a critical point
Here, time $t\leq 0$, $\tau$ is the ramp time, and $r>0$ is the ramp exponent. The power law (1) is assumed only near the end of the ramp, close to the critical point where $\epsilon$ is small. Before the ramp begins at a large initial $\epsilon$, the system is initialized in an easy-to-prepare ground state of the initial Hamiltonian. Thanks to its large energy gap, the initial stage of the evolution is adiabatic. The adiabaticity must fail when the critical point is approached and the gap shrinks to zero. For a long enough ramp time $\tau$ the failure happens close enough to the critical point for the asymptote (1) to be accurate and, furthermore, for the gap to scale as $\Delta\propto|\epsilon|^{z\nu}$, where $z$ and $\nu$ are the dynamical and correlation length exponents, respectively. According to the non-linear KZM [18, 20, 19], see also the review [9], the failure happens when the rate of the transition, $|\dot{\epsilon}/\epsilon|\propto 1/|t|$, becomes comparable to the gap near $-\hat{t}$, where
@@
As discussed at length in Ref. 37, the adiabatic-impulse approximation is not quantitatively accurate, as the correlation range can more than double between $-\hat{t}$ and 00. However, it correctly predicts the power law (3) satisfied by the KZ length $\hat{\xi}$, which is the unique scale of length characterizing the state in the framework of the KZ scaling hypothesis [82, 83, 84]. In particular, the density of point-like excitations scales as
@@
In order to avoid these traps we employ a simple symmetric ramp profile with a time-like parameter $u$:
@@ III An inhomogeneous ramp to a critical point
We consider inhomogeneous ramps that employ the envelope function (10) with a local time-like parameter
@@ IV Quantum Ising chain
Fig. 2(a) shows the final density of excited quasiparticles in a finite chain with open boundary conditions after an inhomogeneous ramp, as a function of the ramp velocity. For smooth ramps with slope $\alpha\ll 1$, we can clearly see suppression of excitations when the ramp is slower than the speed of sound $c=2$. The picture is further corroborated in Fig. 2(b), where the density is scaled by a factor $\alpha^{-2/3}$ and shown as a function of scaled velocity $v/c$. For the smooth ramps, the scaled plots collapse. Setting $s_{c}=0$, the profile (13) becomes $u=\alpha(vt-s)$ and, from the point of view of site $s$, it appears locally like a uniform ramp with a ramp time
@@
As we elaborate in App. A, for supersonic $v>c$, the ramp indeed proceeds as if it were locally uniform, leaving behind the quasiparticle density scaling like $d\propto\tau_{local}^{-2/3}\propto\alpha^{2/3}$. This effective locality follows from the inability to communicate what happened behind the supersonic front to the area ahead of it. However, the effective local time undergoes a dilatation that decorates the density with an extra quasi-Lorentz factor:
@@
Here, $c_{s}$ is a fermionic annihilation operator, and for a given $s$, the sum runs over its two NN sites $s^{\prime}=s+\hat{e}_{x}$ and $s^{\prime}=s+\hat{e}_{y}$. The frequency spectrum of fermionic Bogolubov quasiparticles is
@@
This Fermi point results from merging, as $\lambda\to 1^{-}$, of two Fermi points that coexist within the gapless phase, $10<\lambda<$ 1, located at ±$\pm(k_{F},-k_{F})/\sqrt{2}$, where $k_{F}=\sqrt{2}\arccos\lambda$. Near each of them, we can parameterize $(k_{x},k_{y})=(\pm k_{F}+q+p,\mp k_{F}-q+p)/\sqrt{2}$ and expand the dispersion for small $p,q$. Close to both Fermi points, $\omega^{2}\approx c^{2}_{p}(\lambda)p^{2}+c^{2}_{q}(\lambda)q^{2}$, where $c_{p}$ and $c_{q}$ are the quasiparticle velocities in orthogonal directions. $c_{p}(\lambda)$ monotonically increases with $\lambda$ from $c_{p}(0)=0$, while $c_{q}(\lambda)$ decreases to $c_{q}(1)=0$ at the phase boundary. At $\lambda=1$, the two Fermi points merge at $(k_{x},k_{y})=(0,0)$. Near this double Fermi point, there is a finite speed
@@ V.1 Uniform ramp
In the second line, we kept only the leading term for large $\tau$. Here, the non-universal constants $a_{1,2}$ are coefficients in the asymptotes $\hat{k}_{1}=a_{1}\tau^{-2/3}$ and $\hat{k}_{2}=a_{2}\tau^{-1/3}$. Given that $|p|$ and $|q|$ are excited in the range from 00 up to $\hat{k}_{1}$ and $\hat{k}_{2}$, respectively, equation (27) implies average energy of excited quasiparticles $\propto\tau^{-2/3}$ and the excitation energy density scaling as
@@
For large enough $L$, Figs. 3(a,b) demonstrate the crossover to the adiabatic regime. In a finite system of linear size $L$, the minimal quasimomenta $p,q\approx 1/L$ and, for large enough $L$, it is the nonzero speed $c_{p}=2\sqrt{2}$ in the direction $z=1$ that determines the minimal gap $\propto L^{-1}$ in (26). Accordingly, the KZ scale of length that is relevant for the minimal gap is $\hat{\xi}_{1}\propto\hat{k}_{1}^{-1}\propto\tau^{2/3}$, and the uniform ramp is adiabatic when $\hat{\xi}_{1}\gg L$ or, equivalently, $\tau\gg L^{3/2}$. This prediction agrees with the insets in Figs. 3(a,b) that show, respectively, scaled $L^{3/2}d$ and $L^{5/2}{\rm e}$ as functions of scaled time $L^{-3/2}\tau_{total}$. The collapse of the scaled plots confirms that $\tau\propto L^{3/2}$ marks the crossover to the adiabatic regime.
@@ VI Kitaev model
Here, the sum runs over all fermionic quasiparticle modes of the final Hamiltonian, $p_{j}$ is the excitation probability of the $j$-th mode, and the approximation is accurate when $p_{j}\ll 1$ for all modes. Therefore, $f$ can be thought of as a substitute for the density of quasiparticle excitations $d$ in the case of a nonintegrable system.
Our simulation results are shown in Fig. 9. The exponential growth of the matrix product state’s bond dimension with system size limits the simulations to a $10\times 10$ lattice with open boundary conditions. Fig. 9(a) shows the infidelity as a function of the ramp velocity. It is clearly suppressed at low velocities but the speed of sound is obscured by finite size effects. Nevertheless, as shown in Fig. 9(b), there is a clear advantage of a suitable inhomogeneous ramp over the homogeneous one in terms of the infidelity as a function of the total preparation time. Due to the limited size, $\alpha=1,1/2$ are the optimal slopes as they not only minimize the excitation but also their small width limits the time offset.
@@ Appendix A A supersonic ramp in the 1D quantum Ising model
(88) can be brought to a non-linear Landau-Zener form [90]
@@
Here, $z=4|\tilde{k}|\tilde{t}$ is a new time-like variable and $\delta=4^{-1}\epsilon_{0}^{1/r}|\tilde{k}|^{-(r+1)/r}\tilde{\tau}^{-1}$ is a transition rate. As $\delta$ is the only parameter, the ramp ends at $z=0$, precisely at the anti-crossing, with a transition probability $p_{r}(\delta)$. The probability does not depend on the sign of $\tilde{k}$. It is non-negligible when the transition rate $\delta\gg 1$ or, equivalently, for $|\tilde{k}|$ up to approximately $\tilde{\tau}^{-r/(1+r)}$. We note that for $v\to c^{+}$, the rate $\delta$ shrinks to zero, suggesting adiabaticity for subsonic ramps. In order to obtain the number of excited quasiparticles for $v>c$, the probability $p_{r}(\delta)$ has to be integrated over quasiparticle modes with a proper measure that can be determined by initial conditions.
@@
where $\tilde{k}=k-\frac{\omega_{k}}{v}$. For a sufficiently large initial $\epsilon_{i}$, the quasiparticle dispersion $\omega_{k}$ is nearly flat, and $\tilde{k}$ is just shifted by a constant with respect to $k$. Therefore, the density of excited quasiparticles can be approximated by an integral over $\tilde{k}$:
$d(v>c)$$\approx$ $\int\frac{d\tilde{k}}{2\pi}\leavevmode\nobreak\ p_{r}\left(\delta\right)\leavevmode\nobreak\ =A_{r}\tilde{\tau}^{-r/(1+r)}=$ (96)
@@
Here, $A_{r}$ is a non-universal constant. Numerical integration yields $A_{2}=0.039$, while $A_{2}=0.029$ is the best fit to data obtained by simulation of a supersonic inhomogeneous ramp of the transverse field $g_{s}$. We note that the values of $A_{2}$ provided here are for linear ramp of the transverse field considered in the Appendix, and the values of $A_{2}$ appearing in the main text are for a smooth ramp of both transverse field and couplings considered there. Equation (96) is used in Fig. 2 as a theoretical function with $A_{2}$ being its only fitting parameter. Finally, we notice that for $v\gg c$, the density

### v2

# Inhomogeneous adiabatic preparation of a quantum critical ground state in two dimensions

@@ I Introduction
Figure 1: Inhomogeneous ramp. Here, $\epsilon$ is a dimensionless parameter in a Hamiltonian that measures the distance to the quantum critical point. (a) shows an inhomogeneous ramp in a 1D chain at two different times. The transition begins at the center of the chain, where the critical $\epsilon=0$ is reached first, and from where it expands with velocity $v$ towards the chain’s ends. (b) shows an inhomogeneous ramp in a $10\times 10$ square lattice. The critical region with $\epsilon=0$ expands with velocity $v$ from the center towards the edges of the lattice. Both (a) and (b) demonstrate ramps with slope $\alpha=1$.
When the critical $H_{f}$ is approached uniformly in space, the AP ramp is a “half-KZM” ramp, ending at the critical point instead of crossing it, and the KZM power laws still apply, as discussed at length in the following Sec. II. A full KZM ramp can be made adiabatic by making it inhomogeneous in space [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79]. Initially, the parameter $\lambda(t,{\bf r})$ can be driven across the critical point at the center of the lattice, then the central region with $H_{f}$ is expanded towards the lattice edges with a velocity $v$ slower than the speed of sound at the critical point, see Fig. 1.
An inhomogeneous ramp has also been considered as a means to prepare a critical ground state, either in a 1D chain [80] or a quasi-1D stripe [81]. These studies focused on slightly supersonic ramps that create excitations, which follow the ramp towards the ends of the chain, leaving behind the critical ground state in the bulk. Since the excitations eventually bounce back from the ends, we prefer to focus on subsonic ramps in this work. Furthermore, instead of using excitation energy as a cost function—which under-represents low-frequency excitations that can distort the target state without significantly increasing its energy—we focus on the density of excited quasiparticles (for integrable systems) or, more generally, on the (in)fidelity between the critical ground state and the final state (for nonintegrable systems).
@@ II A uniform ramp to a critical point
Here, time $t\leq 0$, $\tau$ is the ramp time, and $r>0$ is the ramp exponent. The power law (1) is assumed only near the end of the ramp, close to the critical point where $\epsilon$ is small. Before the ramp begins at a large initial $\epsilon$, the system is initialized in an easy-to-prepare ground state of the initial Hamiltonian. Thanks to its large energy gap, the initial stage of the evolution is adiabatic. The adiabaticity must fail when the critical point is approached and the gap shrinks to zero. For a long enough ramp time $\tau$ the failure happens close enough to the critical point for the asymptote (1) to be accurate and, furthermore, for the gap to scale as $\Delta\propto|\epsilon|^{z\nu}$, where $z$ and $\nu$ are the dynamical and correlation length exponents, respectively. According to the nonlinear KZM [18, 20, 19], see also the review [9], the failure happens when the rate of the transition, $|\dot{\epsilon}/\epsilon|\propto 1/|t|$, becomes comparable to the gap near $-\hat{t}$, where
@@
As discussed at length in Ref. 37, the adiabatic-impulse approximation is not quantitatively accurate, as the correlation range can more than double between $-\hat{t}$ and 00. However, it correctly predicts the power law (3) satisfied by the KZ length $\hat{\xi}$, which is the unique scale of length characterizing the state in the framework of the KZ scaling hypothesis [82, 83, 84]. In particular, the density of pointlike excitations scales as
@@
In order to avoid these traps we employ a simple symmetric ramp profile with a timelike parameter $u$:
@@ III An inhomogeneous ramp to a critical point
We consider inhomogeneous ramps that employ the envelope function (10) with a local timelike parameter
@@ IV Quantum Ising chain
Figure 2(a) shows the final density of excited quasiparticles in a finite chain with open boundary conditions after an inhomogeneous ramp, as a function of the ramp velocity. For smooth ramps with slope $\alpha\ll 1$, we can clearly see suppression of excitations when the ramp is slower than the speed of sound $c=2$. The picture is further corroborated in Fig. 2(b), where the density is scaled by a factor $\alpha^{-2/3}$ and shown as a function of scaled velocity $v/c$. For the smooth ramps, the scaled plots collapse. Setting $s_{c}=0$, the profile (13) becomes $u=\alpha(vt-s)$ and, from the point of view of site $s$, it appears locally like a uniform ramp with a ramp time
@@
As we elaborate in Appendix A, for supersonic $v>c$, the ramp indeed proceeds as if it were locally uniform, leaving behind the quasiparticle density scaling like $d\propto\tau_{local}^{-2/3}\propto\alpha^{2/3}$. This effective locality follows from the inability to communicate what happened behind the supersonic front to the area ahead of it. However, the effective local time undergoes a dilatation that decorates the density with an extra quasi-Lorentz factor:
@@
Here, $c_{s}$ is a fermionic annihilation operator, and for a given $s$, the sum runs over its two NN sites $s^{\prime}=s+\hat{e}_{x}$ and $s^{\prime}=s+\hat{e}_{y}$. The frequency spectrum of fermionic Bogoliubov quasiparticles is
@@
This Fermi point results from merging, as $\lambda\to 1^{-}$, of two Fermi points that coexist within the gapless phase, $10<\lambda<$ 1, located at ±$\pm(k_{F},-k_{F})/\sqrt{2}$, where $k_{F}=\sqrt{2}\arccos\lambda$. Near each of them, we can parametrize $(k_{x},k_{y})=(\pm k_{F}+q+p,\mp k_{F}-q+p)/\sqrt{2}$ and expand the dispersion for small $p,q$. Close to both Fermi points, $\omega^{2}\approx c^{2}_{p}(\lambda)p^{2}+c^{2}_{q}(\lambda)q^{2}$, where $c_{p}$ and $c_{q}$ are the quasiparticle velocities in orthogonal directions. $c_{p}(\lambda)$ monotonically increases with $\lambda$ from $c_{p}(0)=0$, while $c_{q}(\lambda)$ decreases to $c_{q}(1)=0$ at the phase boundary. At $\lambda=1$, the two Fermi points merge at $(k_{x},k_{y})=(0,0)$. Near this double Fermi point, there is a finite speed
@@ V.1 Uniform ramp
In the second line, we kept only the leading term for large $\tau$. Here, the non-universal constants $a_{1,2}$ are coefficients in the asymptotes $\hat{k}_{1}=a_{1}\tau^{-2/3}$ and $\hat{k}_{2}=a_{2}\tau^{-1/3}$. Given that $|p|$ and $|q|$ are excited in the range from 00 up to $\hat{k}_{1}$ and $\hat{k}_{2}$, respectively, Eq. (27) implies average energy of excited quasiparticles $\propto\tau^{-2/3}$ and the excitation energy density scaling as
@@
For large enough $L$, Figs. 3(a) and  3(b) demonstrate the crossover to the adiabatic regime. In a finite system of linear size $L$, the minimal quasimomenta $p,q\approx 1/L$ and, for large enough $L$, it is the nonzero speed $c_{p}=2\sqrt{2}$ in the direction $z=1$ that determines the minimal gap $\propto L^{-1}$ in (26). Accordingly, the KZ scale of length that is relevant for the minimal gap is $\hat{\xi}_{1}\propto\hat{k}_{1}^{-1}\propto\tau^{2/3}$, and the uniform ramp is adiabatic when $\hat{\xi}_{1}\gg L$ or, equivalently, $\tau\gg L^{3/2}$. This prediction agrees with the insets in Figs. 3(a) and 3(b) that show, respectively, scaled $L^{3/2}d$ and $L^{5/2}{\rm e}$ as functions of scaled time $L^{-3/2}\tau_{total}$. The collapse of the scaled plots confirms that $\tau\propto L^{3/2}$ marks the crossover to the adiabatic regime.
@@ VI Kitaev model
Here, the sum runs over all fermionic quasiparticle modes of the final Hamiltonian, $p_{j}$ is the excitation probability of the $j$-th mode, and the approximation is accurate when $p_{j}\ll 1$ for all modes. Therefore $f$ can be thought of as a substitute for the density of quasiparticle excitations $d$ in the case of a nonintegrable system.
Our simulation results are shown in Fig. 9. The exponential growth of the matrix product state’s bond dimension with system size limits the simulations to a $10\times 10$ lattice with open boundary conditions. Figure 9(a) shows the infidelity as a function of the ramp velocity. It is clearly suppressed at low velocities but the speed of sound is obscured by finite size effects. Nevertheless, as shown in Fig. 9(b), there is a clear advantage of a suitable inhomogeneous ramp over the homogeneous one in terms of the infidelity as a function of the total preparation time. Due to the limited size, $\alpha=1,1/2$ are the optimal slopes as they not only minimize the excitation but also their small width limits the time offset.
@@ Appendix A A supersonic ramp in the 1D quantum Ising model
(88) can be brought to a nonlinear Landau-Zener form [90]
@@
Here, $z=4|\tilde{k}|\tilde{t}$ is a new timelike variable and $\delta=4^{-1}\epsilon_{0}^{1/r}|\tilde{k}|^{-(r+1)/r}\tilde{\tau}^{-1}$ is a transition rate. As $\delta$ is the only parameter, the ramp ends at $z=0$, precisely at the anti-crossing, with a transition probability $p_{r}(\delta)$. The probability does not depend on the sign of $\tilde{k}$. It is non-negligible when the transition rate $\delta\gg 1$ or, equivalently, for $|\tilde{k}|$ up to approximately $\tilde{\tau}^{-r/(1+r)}$. We note that for $v\to c^{+}$, the rate $\delta$ shrinks to zero, suggesting adiabaticity for subsonic ramps. In order to obtain the number of excited quasiparticles for $v>c$, the probability $p_{r}(\delta)$ has to be integrated over quasiparticle modes with a proper measure that can be determined by initial conditions.
@@
where $\tilde{k}=k-\frac{\omega_{k}}{v}$. For a sufficiently large initial $\epsilon_{i}$, the quasiparticle dispersion $\omega_{k}$ is nearly flat, and $\tilde{k}$ is just shifted by a constant with respect to $k$. Therefore the density of excited quasiparticles can be approximated by an integral over $\tilde{k}$:
$d(v>c)$$\approx$ $\int\frac{d\tilde{k}}{2\pi}\leavevmode\nobreak\ p_{r}\left(\delta\right)\leavevmode\nobreak\ =A_{r}\tilde{\tau}^{-r/(1+r)}$ (96)
@@
Here, $A_{r}$ is a nonuniversal constant. Numerical integration yields $A_{2}=0.039$, while $A_{2}=0.029$ is the best fit to data obtained by simulation of a supersonic inhomogeneous ramp of the transverse field $g_{s}$. We note that the values of $A_{2}$ provided here are for linear ramp of the transverse field considered in the Appendix, and the values of $A_{2}$ appearing in the main text are for a smooth ramp of both transverse field and couplings considered there. Equation (96) is used in Fig. 2 as a theoretical function with $A_{2}$ being its only fitting parameter. Finally, we notice that for $v\gg c$, the density